OpenStudy (richley):

I need some help with Sigma Notation! I'm in Honors PreCalculus and am confused!

OpenStudy (richley):

Obtain a formula for\[\sum_{k=1}^{n}k^m\]

OpenStudy (richley):

I seem to be having lots of trouble with Sigma Notation. It's starting to get frustrating because I have 3 problems for my project that I can't quite solve and it's due tomorrow.

OpenStudy (eliesaab):

Are you sure you typed your question right?

OpenStudy (richley):

Yes I'm sure. It's k^m and not k^n. That would make it much simpler.

OpenStudy (eliesaab):

Does the sum start from k=1 or m=1?

OpenStudy (richley):

k=1

OpenStudy (eliesaab):

In this case, this is not a easy sum.

OpenStudy (richley):

If it helps at all, it's given as a tip that \[\sum_{k-1}^{n}k^2 = \frac{ n(n+1)(2n+1) }{ 6}\]

OpenStudy (richley):

but I don't quite see how that helps

OpenStudy (richley):

Well it never states m as equal to n+1

OpenStudy (richley):

Why would that be the case?

OpenStudy (sweetburger):

Sorry i keep having typos one sec

OpenStudy (eliesaab):

\[ \sum _{k=1}^n k^3=\frac{1}{4} n^2 (n+1)^2 \]

OpenStudy (richley):

Ah, so how would that relate to k^m? It's weird since m is not defined in the question

OpenStudy (eliesaab):

\[ \sum _{k=1}^n k^4=\frac{1}{30} n (n+1) (2 n+1) \left(3 n^2+3 n-1\right) \] and so fourth, but there is no easy formula for any m

OpenStudy (richley):

That makes this really tricky doesn't it

OpenStudy (richley):

I was able to go as far as to derive that

OpenStudy (richley):

But I really can't figure out how to do it for the mth power, since the pattern isn't really that easy to determine

OpenStudy (eliesaab):

For a general m, the sum is defined in terms of waht is called the Harmonic Number

OpenStudy (richley):

can that be used to generalize a formula here?

OpenStudy (richley):

We haven't learned about harmonic numbers yet i dont think

OpenStudy (richley):

I can give you a summary of previous knowledge of this unit?

OpenStudy (richley):

1) Matrices 2) Geometric Sequences and Series 3) Arithmetic Sequences and Series 4) Pascal's Triangle 5) The Binomial Theorem

OpenStudy (richley):

That still uses the Harmonic Number

OpenStudy (richley):

Is there a way to do it without using that to generalize?

OpenStudy (eliesaab):

This is the only way. That is why I asked if the summation for k=1, or m=1. In case, it is m=1, to n. It is a geometric series.

OpenStudy (richley):

This is a really confusing problem isn't it?

OpenStudy (richley):

The next one is a bit simpler, maybe if you walk me through it i'll be able to understand how this notation works with exponents?

OpenStudy (richley):

Find a formula for \[\sum_{k=1}^{n}k^3\]

OpenStudy (eliesaab):

This summation is done in Calculus for specific m. Like m=1, m=2, m=3 but never for a general m

OpenStudy (eliesaab):

That is easy, I gave you the answer above

OpenStudy (richley):

Yes but how is that answer obtained?

OpenStudy (richley):

I don't really need the answers exactly, I need to know how to obtain those answers.

OpenStudy (eliesaab):

Visit the above link

OpenStudy (richley):

Will you be here in a few minutes? I wanna read through that and work it out.

OpenStudy (eliesaab):

There are several approaches. All of them lead to the solution.

OpenStudy (richley):

What's the link to the solution for the k^4 one?

OpenStudy (richley):

Ah! I think I got the formula for the first one

OpenStudy (richley):

Thanks a ton for your help. I believe the answer is \[\sum_{i=1}^{n}k^{m-1}\sum_{k=1}^{n}1\]